OpenStudy (anonymous):

A Piper Cub has an average air speed that is 10mph faster than a Cessna 150 airplane. If the combined distance traveled by these two small planes is 690 miles after 3 hr, what is the average speed of each plane?

OpenStudy (anonymous):

let x be speed 1, y be speed 2. 3(x+y) = 690. x = y + 10. So Y=110 X= 120 is this correct?

OpenStudy (anonymous):

v1 = 10 + v2

OpenStudy (anonymous):

so I'm wrong?

OpenStudy (anonymous):

well im doing it differently

OpenStudy (anonymous):

do you agree that v1 = 10 + v2

OpenStudy (anonymous):

v1 is the speed of the piper cub

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok the distance we know is d = v* t , where v is the speed or rate, like d = rt

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the total distance you add the distances of the two planes , so v1*3 + (10+v2) *3 = 690

OpenStudy (anonymous):

so you have a system of equations ,

OpenStudy (anonymous):

since you know that v1 = 10 + v2, you can substitute that into the second equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(10 + v2) *3 + (10+v2)*3 = 690

OpenStudy (anonymous):

maybe someone should check my work

OpenStudy (anonymous):

\[\sqrt{105}\]

OpenStudy (anonymous):

v =10.25695

OpenStudy (anonymous):

i got v2 = 105

OpenStudy (anonymous):

as did I

OpenStudy (anonymous):

so v1 = 115

OpenStudy (anonymous):

hmm, so The first one I did was wrong

OpenStudy (anonymous):

no no nio

OpenStudy (anonymous):

i made a mistake

OpenStudy (anonymous):

we had v1 *3 + v2 *3 = 690

OpenStudy (anonymous):

3 ( v1 +v2 ) = 690

OpenStudy (anonymous):

substitute v1 = 10 + v2

OpenStudy (anonymous):

3 ( 10 + v2 + v2) = 690

OpenStudy (anonymous):

so v2 = 110

OpenStudy (anonymous):

so v1 = 100

OpenStudy (anonymous):

i meant

OpenStudy (anonymous):

oh, I got 120, let me check my math

OpenStudy (anonymous):

v1 = 120

OpenStudy (anonymous):

you did everything right

OpenStudy (anonymous):

AWESOME! thank you sooo much!